OpenStudy (anonymous):

how do I solve the equation x^2+2x+8=0 and write the complex solution in standard form

OpenStudy (anonymous):

Use the quadratic formula. \[ x=\frac{-b\pm i \sqrt{4ac- b^2}}{2a}, \text { if } 4ac - b^2>0 \]

OpenStudy (anonymous):

a=1, b=2, c=8 \[ 4ac - b^2= 4(1)(8) - (2)^2= 32 -4= 28 \]

OpenStudy (anonymous):

\[ x=\frac { -2 \pm i \sqrt{28}}{2}=-1 \pm i \frac{2 \sqrt 7}2\\ x_1 = -1 + \sqrt 7\ i \\ x_2 = -1 - \sqrt 7\ i \\ \]